OpenStudy (anonymous):

Which is the solution to |2x - 4| > 8

sam (.sam.):

I'd square both sides to solve (2x-4)^2>8^2

sam (.sam.):

Do you know how to continue from here?

OpenStudy (anonymous):

I did it but I end up getting the square root of x = 24 and then when I factor it , the answer is not one of the choices

OpenStudy (compassionate):

Hello, You are solving absolute values. The first step is to eliminate the 4. Do you know how?

OpenStudy (anonymous):

don't you just add 4 to the other side

OpenStudy (compassionate):

Correct. Absolute values means that you are 4 spaces from zero. So, add the 4 to the eight.

OpenStudy (anonymous):

which gives me 2x>12

OpenStudy (compassionate):

Now divide the 2.

OpenStudy (anonymous):

X > 6 , but that is not one of my choices that's why I have been confused

OpenStudy (compassionate):

Okay, I understand what you're asking. One second please.

OpenStudy (anonymous):

x = 2 or x = 6 -2 < x < 6 x < -2 or x > 6 x < -6 or x > 2

OpenStudy (compassionate):

\[x > 6 ----- x < -2\]

OpenStudy (compassionate):

The anwer would be C. I'll allow Mathslover to explain how we came to the answer. I see him typing.

mathslover (mathslover):

well please wait then for 2 minutes.

OpenStudy (compassionate):

If you don't feel like explain, then I can.

mathslover (mathslover):

see we have : \(\large |2x-4| > 8\) Now I can write this as : \(\large 2|x-2| > 2 * 4\) \(\implies |x-2| > 4\) So we have two cases here : either x-2 > 4 or x-2 < -4 so : \(\large x > 6\) as x-2 > 4 and x < -2 as x-2 < -4 so I have : \(\large x \in (-\infty , -2 ) \space \cap (6 , \infty ) \)

OpenStudy (anonymous):

ohhh because its absolute value you have to do the both positive and negative forms... thanks everybody for helping I really appreciate it

mathslover (mathslover):

This is the shortest and logical method for solving this question.

OpenStudy (compassionate):

Yes - Brand, because the absolute value of any number means the numbers that are x spaces from zero. So if I had I 4 I, I positive and negative 4 are four spaces from zero.

mathslover (mathslover):

The method suggested by @.Sam. was also good but that would have been longer than the above mentioned method. And thanks @Compassionate for giving me the chance to put here my opinion .

OpenStudy (compassionate):

No problem. I hate to snipe other users.

mathslover (mathslover):

:)

OpenStudy (compassionate):

For some addition methods, the way I solved it was: |2x-4| > 8 Either: 2x - 4 > 8 2x > 12 x > 6 Or: -|2x - 4| > 8 -2x + 4 > 8 -2x > 4

OpenStudy (compassionate):

x > -2

mathslover (mathslover):

And , \(\huge{\color{orange}{\textbf{Welcome}} \space \color{blue}{\mathbb{To}} \space \color{purple}{\cal{OpenStudy}}}\) @brandistephens . Keep asking and answering questions here on OpenStudy. Here is an awesome openstudy guide for the new users like you : http://prezi.com/fs3hqdpcopic/an-unofficial-guide-to-openstudy/ Best of luck!